Let p(x)=ax3−x, where a is constant and a>0.

Find the local maxima and minima of p.
(Enter your maxima and minima as comma-separated xvalue,classification pairs. For example, if you found that x=−2 was a local minimum and x=3 was a local maximum, you should enter (-2,min), (3,max). If there were no maximum, you must drop the parentheses and enter -2,min.)

maxima and minima:
1
Expert's answer
2020-03-16T11:46:42-0400


Apply the First Sufficient Condition For The Extremum. Find the first derivative.

"p'(x) = 3ax^2-1"

Find x0 where the first derivative is equal to zero.

"3ax^2-1=0"

"x^2=x_1=\frac{1}{3a}"

"x_1=\sqrt{\frac{1}{3a}}" and "x_2=-\sqrt{\frac{1}{3a}}"

Check that "p'(x)" changes the sign in x1 and x2 . "p'(x)" is a parabola with vertex "x_0=\frac{-b} {2a}=0; y_0=p'(x_0)=-1"

It means that parabola branches cross the Ox in x1 and x2 . 

The graph of derivative is shown on Picture 1:



Picture 1

As we can see the vertex is in 3 and 4 quarters where the function has minus sign and after crossing the Ox in x1 and x2 function goes to 1 and 2 quarters and changes the sign to plus.

It means that extrema are found.

Find the value of function p(x) in extrema and compare.

"p(x_1)=ax_1^3-x_1=-\frac{2}{3}*\sqrt{\frac{1}{3a}}"

"p(x_2)=\frac{2}{3}*\sqrt{\frac{1}{3a}}"

As "p(x_2)>p(x_1)" then x2 is a point of maximum and x1 is a point of minimum.


Answer

"(\sqrt{\frac{1}{3a}},min)" 

"(-\sqrt{\frac{1}{3a}},max)"